OpenStudy (anonymous):

How do i solve this? 0.0567kg / 2.364*10^2mL=

OpenStudy (abhisar):

Hello @lamp12345 !

OpenStudy (abhisar):

\(\huge \frac{Mass}{Volume}=Density\\ \Rightarrow \huge \frac{5.67\times 10^{-2}}{2.364\times 10^2}=2.39\times 10^{-4}Kg/mL\)

OpenStudy (zale101):

@Abhisar divided the decimal parts together, \[5.67/2.364≈2.39\](round to the hundredth place) for the exponential part, he subtracted the exponents \[10^{−2−2}=10^{−4}\]\[2.39×10^{−4} Kg/mL\]You subtract the exponents if you were dealing with division of scientific notation or any algebraic expression with exponents in them. With multiplication of scientific notations or any algebraic expressions, you add the exponents :)